# FRAMES: Boosting LLMs with A Four-Quadrant Multi-Stage Pretraining Strategy

### v1

@@ B.6 Smoother LOSS Curve from FRAMES
Our analysis reveals that the loss curve from FRAMES has the lowest high-frequency energy proportion at 0.02%, significantly lower than that of PDPC and Random. This suggests that FRAMES allows the model to converge more stably, and reduces the impact of gradient fluctuations during training. This improvement is due to our four-stage training strategy, which organizes similar data into consecutive batches, thus stabilizing the model training process.
B.7 Ablation Studies
@@
Figure 16: PPL related ablation study (2 quadrants)
@@
Figure 17: PPL related ablation study (3 quadrants)
@@
Figure 18: PD related ablation study (2 quadrants)
@@
Figure 19: PD related ablation study (3 quadrants)
Appendix C Data Cases
Table LABEL:PD_quantiles presents samples extracted from each quadrant. The data in the $Q_{1}$ quadrant primarily consists of samples with low PPL and low PD. These samples exhibit homogeneity and a high degree of duplication. The $Q_{2}$ quadrant mainly contains data with low PPL but high PD, featuring text expressions that conform to standard written language conventions, with samples being relatively longer and more complex. The data in $Q_{3}$ and $Q_{4}$, with high PPL and a mix of low and high PD, often feature creative, informal, or unexpected language that doesn’t follow typical grammar rules or standard phrasing, differing significantly from usual writing or speaking styles.
@@ Appendix C Data Cases
Sample 1: …The need to practice good self-care doesn’t change in this working environment, but how you accomplish this goal might. Much of Arel’s own self-care regimen needed to be adjusted."I was used to weekly massage and monthly chiropractic care. That was gone," she explains. "I am used to runs and yoga and time to meditate in complete silence. That was gone, too." And although there was an onboard fitness center available, Arel often found that she couldn’t get a workout in during the day and was too tired at the end of the day to be invested in the results …
Sample 2: … I have to ask you, why’d you–wha–wha–why are you peeing right here?Creepy Guy: What?Kumar: I mean… why’d you pee right next to me when you could like, choose that bush, or–?Creepy Guy: Well, this bush looked like I should pee on it. Why are you peeing on it?Kumar: Well, no one was here when I chose this bush.Creepy Guy: Oh, so you get to pee on it and no one else does? Huh?Kumar: No, it’s just–I just–Creepy Guy: This your bush? You have a special bond with this bush?Kumar: No, I just thought that–Creepy Guy: You the king of the forest…
@@
Sample 1: … boolean insertventas() String sql "INSERT INTO ventas (id_venta, venFechaventa, venId_cliente, venIdadministrador, venTotalventa) VALUES (NULL, ’" + vent() "’, ’" clasu.getId_usuarios() "’, ’1’, ’" + pnlProductos.totall + "’)"; try con cn.getConnection(); ps = con.prepareStatement(sql); ps.executeUpdate(); return true; catch (SQLException ex) Logger.getLogger(LogicaSql.class…
Sample 2: … In terms of providing shorter stay parking, Bell Street multi storey car park is identified as a long stay car park, and the tariffs are so designed to encourage the use of the facility by all day / half day parkers with more flexible tariffs available at other car parks and the on street spaces around the vicinity allow for parking for up to one hour.I have commented that there is no short term (30 minutes to 2 hour) parking available at the West Bell Street multi-storey car park and the …
@@
Sample 1: … Maybe it just sagslike a heavy load. Or does it explode?   by Langston HughesIn 1849, Elizabeth Blackwell became the first woman to graduate from a U.S. medical school in N.Y.In 1864, Rebecca Lee Crumpler became the first black woman to graduate from a U.S. medical school in New England.In 1915, women represented approximately 5% of the physician workforce in the U.S.In 1983, women represented approximately 1/3 of U.S. medical school matriculants.In 1976, Grady doctor and cardiologist Nanette K. Wenger is recognized as one of Time Magazine’s Women of the year, and in 1998 she is named Chief of Cardiology.In 2006, women represented nearly one half of all U.S. medical …
Sample 2: … FILED UNDER SEAL PURSUANT TO PROTECTIVE ORDER rise to a direct infringement claim against it. See, e.g., Akami Techs., Inc. v. Limelight Networks, Inc., 797 F.3d 1020, 1023 (Fed. Cir. 2015) (noting entities are liable for performance they control). The evidence further shows that Badoo Software Limited and Badoo Limited are also intimately involved in Badoo Trading’s creation and ownership of the infringing Bumble application. Ex. A at Recital (c). Badoo Limited’s records indicate that its "principal activity" is "to provide mobile development services to Badoo Software Limited, another entity in the Worldwide Vision group…
@@
Sample 1: … I myself should be a castaway.Young’s Literal: but I chastise my body, and bring it into servitude, lest by any means, having preached to others – I myself may become disapproved.As noted earlier, Paul now applies the example from the Greek sports arena directly to himself ("I discipline… I myself") and does so that he might present himself as an example or model for other believers to imitate (cp 1Co 4:16, 11:1, 1Th 1:6, cp Heb 6:12, He 13:7, 3Jn 1:11)…
Sample 2: … (B) of from about 0.1 to about 10.0% w/v of a bioadhesive polymeric stabilizer selected from the group consisting of:(i) polyethylene-polypropylene oxide tri-block co-polymers of the formula;(polyethylene oxide)a -(polypropylene oxide)b -(polyethylene oxide)c wherein PA4 a is 46, 52, 62, 75, 97, 98, 122, or 128; PA4 b is 16, 30, 35, 39, 47, 54, or 67; and PA4 c is 46, 52, 62, 75, 97, 98, 122, or 128;(ii) polyvinyl alcohol,(iii) polyvinyl pyrr…
@@
Appendix D Prompts for Case Study
Prompts for Property Recognition You are a language model training data annotator. Your task is to identify whether the given text possesses the following characteristic: {Property} The text to be annotated is: {text} Please determine whether the given text possesses this characteristic according to the above rules. The output format should be "Because…, my answer is X’." where X must be either "yes" or "no." You should remain objective and refrain from adding any further comments after making your choice. 1. Does the text contain polysemous words? Polysemous words may make understanding more difficult. 2. Does the text use specialized terminology? Specialized terminology may require specific domain knowledge to understand. 3. Does understanding the text require specific cultural background knowledge? Cultural background dependence may increase the complexity of understanding. 4. Does the text require logical reasoning to understand? Logical reasoning adds depth to understanding. 5. Does the text contain elements of humor? Humor may affect the way the text is understood. 6. Does the text explore ethical or moral issues? This may increase the depth of thought. 7. Does the text use complex sentence structures? Complex sentence structures may increase the difficulty of understanding. 8. Does the text contain scientific or technical concepts? These concepts may require specific knowledge to understand. 9. Does the text express obvious emotional tones? Emotional tones may affect the understanding of the text. 10. Does understanding the text require additional background knowledge? Background knowledge requirements may affect the comprehensibility of the text.

### v4

@@ B.7 Smoother LOSS Curve from FRAME
Our analysis reveals that the loss curve from FRAME has the lowest high-frequency energy proportion at 0.02%, significantly lower than that of PDPC and Random. This suggests that FRAME allows the model to converge more stably, and reduces the impact of gradient fluctuations during training. This improvement is due to our four-stage training strategy, which organizes similar data into consecutive batches, thus stabilizing the model training process.
B.8 Ablation Studies
@@
Figure 16: PPL related ablation study (2 quadrants).
@@
Figure 17: PPL related ablation study (3 quadrants).
@@
Figure 18: PD related ablation study (2 quadrants).
@@
Figure 19: PD related ablation study (3 quadrants).
Appendix C Prompts for Case Study
Prompts for Property Recognition You are a language model training data annotator. Your task is to identify whether the given text possesses the following characteristic: {Property} The text to be annotated is: {text} Please determine whether the given text possesses this characteristic according to the above rules. The output format should be "Because…, my answer is X’." where X must be either "yes" or "no." You should remain objective and refrain from adding any further comments after making your choice.
The property comes from the following rules:
  1. 1.
    Does the text contain polysemous words? Polysemous words may make understanding more difficult.
  2. 2.
    Does the text use specialized terminology? Specialized terminology may require specific domain knowledge to understand.
  3. 3.
    Does understanding the text require specific cultural background knowledge? Cultural background dependence may increase the complexity of understanding.
  4. 4.
    Does the text require logical reasoning to understand? Logical reasoning adds depth to understanding.
  5. 5.
    Does the text contain elements of humor? Humor may affect the way the text is understood.
  6. 6.
    Does the text explore ethical or moral issues? This may increase the depth of thought.
  7. 7.
    Does the text use complex sentence structures? Complex sentence structures may increase the difficulty of understanding.
  8. 8.
    Does the text contain scientific or technical concepts? These concepts may require specific knowledge to understand.
  9. 9.
    Does the text express obvious emotional tones? Emotional tones may affect the understanding of the text.
  10. 10.
    Does understanding the text require additional background knowledge? Background knowledge requirements may affect the comprehensibility of the text.
Appendix D Data Cases
Table LABEL:PD_quantiles presents samples extracted from each quadrant.
@@ Appendix D Data Cases
Sample 1: … 112. Ra4+ Kd5 113. Ra5+ Kc4 114. Ra6 Qd7+ 115. Kb6 Kb4 116. Ra7 Qd6+ 117. Kb7 Kb5 118. Kc8 Qf8+ 119. Kb7 Qd8 120. Ra1 Qe7+ 121. Kc8 Qe8+ 122. Kb7 Qe4+ 123. Kc8 Qc4+ 124. Kb7 Qf7+ 125. Kc8 Qg8+ 0-1 Analysis of Utegaliyev - Goh from the Baku Olympiad - Life is tough! ***Addendum: The Rook & Bishop vs Rook ending at move 102 was covered specifically by Karsen Muller & Frank Lamprecht on page 301 of Fundamental Chess Endings. My annotations are slightly confusing there …
Sample 2: … Women With an Alcoholic Parent Have More Risk Factors Ray Kachatorian / Photographer’s Choice / Getty Images There are differences in how parental alcoholism affects daughters as opposed to how it affects sons, particularly when it comes to psychopathology, or mental health disorders, in each gender. Daughters of alcoholics are affected by a parent’s alcoholism in many of the same ways that sons are. Both are at higher risk of developing alcohol abuse disorders compared to children of non-alcoholic parents. But there are some differences in how women are influenced, scientists say …
@@
Sample 1: … pension advice be sought in the case of South Africa. Residents in South Africa are subject to tax (up to 40%) on their total income no matter the source-country. Foreign pensions are exempt from this. Any UK pension, as long as it was not ’received or accrued’ from or in South Africa, and is ’in consideration of past employment’ elsewhere, is exempt from income tax. This specifically excludes anyone whose pension has arisen from time in public office in South Africa. For residents of South Africa, …
Sample 2: … , Bengaluru (As Prepared for Delivery) Honorable President Pranab Mukherjee, Ministers, Excellencies, Distinguished guests: Thank you, it is such a great honor to be here today, accepting this award among such distinguished company. Let me first thank His Excellency the Honorable President Mukherjee, Prime Minister Modi, and the Government of India2̆013 2̆026 By U.S. Mission India | 9 January, 2017 | Topics: Chennai, Press Releases, Speeches | Tags: Bilateral relationship Remarks by Ambassador Richard R. Verma at the inaugural plenary of Indo-Asia Connectivity for Shared Prosperity …
@@
Sample 1: … information may be obtained from CPO. The Anti-Kickback Act of 1986 (41.U.S.C.51-58) was passed to deter subcontractors from making payments and contractors from accepting payments for the purpose of improperly obtaining or rewarding favorable treatment in connection with a prime contract or subcontract. Imposes criminal penalties on any person who knowingly and willfully engages in the prohibited conduct addressed in paragraph (a) of this subsection. Provides for the recovery of civil penalties by the United States from any person who knowingly engages in such prohibited conduct and from any person whose employee, subcontractor, or subcontractor employee provides, accepts, or charges a kickback …
Sample 2: … also wrote Brave New World and numerous essays. Consider some of his words as follows: Ẅe are in a process of developing a whole series of techniques which will enable the controlling oligarchy to get people to actually love their servitude. A really efficient totalitarian state would be one in which the all-powerful executive of political bosses and their army of managers control a population of slaves who do not have to be coerced, because they love their servitude. To make them love it is the task assigned, in present-day totalitarian states, to ministries of propaganda …
@@
Sample 1: … in which 3000 people died in a single atrocity, to one’s horror at the deaths of ten and perhaps sometimes twenty times as many in each of the bombings of such places as Hamburg, Dresden, Tokyo, Hiroshima and Nagasaki. This way of grasping the purport of what area bombing meant, really meant, is vital to making a difference to how we behave and what we accept today in the conduct of conflicts. There is nothing abstract or theoretical about the mass murder in which bombing consists: it is real and terrible, and anything that drives the point home has its place in the debate …
Sample 2: … and China in Africa. Ba’ath Party Dominance and Mistakes. Iraq. A Fragile Country. Burkina Faso. The Church under attack. DR Congo. Insecurity and bad governance contribute to the spread of Ebola. Drug Trafficking. RD.Congo. Beatification of twenty martyr missionaries on track. Africa. Start-ups rolling out. Between Maras and Pandilla El Salvador. A Country In Transition. Nigeria. Troubles in Kano. A New President. The Church of Africa. A Return To Its Origins. Mexico. Üntil dignity becomes the custom.̈ Philippine. Modern Day Missionaries of the World. The Gumuz People, their culture. The Religious Universe of the Gumuz. The Gumuz in front of a changing world. Africa. The art of food. Ancient flavours and genuine ingredients …